On December 31, 2009, Thomas, Inc. borrowed $1,120,000 on a 12%, 15-year mortgage note payable. The note is to be repaid in equal semiannual installments of $81,366 (payable on June 30 and December 31). Prepare journal entries to reflect (a) the issuance of the mortgage note payable, (b) the payment of the first installment on June 30, 2010, and (c) the payment of the second installment on December 31, 2010. Round amounts to the nearest dollar.
Answer #1

Journal

Date Account title Debit Credit
Dec. 31, 2009 Cash 1,120,000
Mortgage note payable 1,120,000
June 30, 2010 Mortgage note payable
Interest expense 67,200
Cash 81,366
Dec. 31, 2010 Mortgage note payable 15,016
Interest expense 66,350
Cash 81,366

Interest expense to be paid on June 30, 2010 = 1,120,000 x 12/100 x 6/12

= $67,200

Hence, principal included in the first installment = 81,366 - 67,200

= $14,166

Interest expense to be paid on Dec. 31 30, 2010 = (1,120,000 - 14,166) x 12/100 x 6/12

= $66,350

Hence, principal included in the second installment = 81,366 - 66,350

= $15,016

Kindly comment if you need further assistance. Thanks